# Non-Unitary Quantum Many-Body Dynamics using the Faber Polynomial Method

### v1

@@ 1 Introduction
Finally, as a last example of non-unitary Markovian dynamics, we can consider the one generated by a purely non-Hermitian Hamiltonian. Non-Hermitian physics emerges intrinsically in various domains of physics beyond quantum physics, encompassing photonics [22], hydrodynamics [23] and active matter [24, 25]. In the context of open quantum systems, a non-Hermitian evolution can be obtained by post-selecting the quantum trajectories corresponding to no quantum jumps, i.e. the no-click limit [26]. Non-Hermitian quantum systems show anomalous static and dynamical properties, which are attracting widespread interest. Among these, we mention the unconventional propagation of quantum correlations [27, 28, 29], the distinct entanglement phase transitions generated by time evolution [30, 31, 32, 33, 34] or their extraordinary sensitivity to boundary conditions, also known as the skin effect [35, 36]. The latter manifests itself through the unusual localisation of all single-particle eigenstates at the system’s edges under open boundary conditions  [37, 38, 39, 40, 41]. Furthermore, it has unique signatures in the dynamics, leading to non-reciprocal transport[42].
@@ 3 Faber Polynomial Method
where we define |$\ket{\Psi_{n}}=\mathcal{P}_{n}\left(\mathcal{H}\right)\ket{\Psi\left(t_{0}\right)}$. As will be demonstrated, the coefficients $c_{n}(\delta t)$ decrease as the order $n$ increases. Moreover, the states |$\ket{\Psi_{n}}$ are efficiently computed through the recurrence relations that the polynomials satisfy. To compute the subsequent level of the expansion, the main computational task is the application of the system’s Hamiltonian onto a particular state. Consequently, the most demanding operation involves only multiplying the Hamiltonian by a limited group of vectors, leading to a resource usage that increases linearly with $\dim\left(\mathcal{H}\right)$ for sparse matrices or quadratically with $\dim\left(\mathcal{H}\right)$ for dense matrices. Linear scaling is expected for Hamiltonians characterising a system with short-range interactions or hoppings. These principles are exactly those underpinning Kernel Polynomial Methods [60], which have become an essential computational resource in the field of condensed matter physics.
@@ 3.2 Non-Unitary Evolution
In order to perform one time step of evolution, one simply has to truncate Eq. 10 up to the desired order $N_{p}$, and calculate the associated states |$\ket{\Psi_{n}}$. Through the relations in Eq. 12, one never has to store the Hamiltonian matrix, only needing to store in memory at most the previous two states, |$\ket{\Psi_{n-1}}$ and |$\ket{\Psi_{n}}$, to compute the following term of the expansion |$\ket{\Psi_{n+1}}$. Furthermore, the expansion coefficients can be computed once at the beginning of the algorithm, as they depend only on the chosen time step. Given this, the computation of the state of the system after a time step scales linearly in the number of polynomials and linearly in the Hilbert space dimension (assuming that the Hamiltonian has a sparse representation on the used basis). This scaling can be improved by using parallelisation techniques and making use of the underlying symmetries of the Hamiltonian [63].
@@ 4.1 Benchmark: Dynamics of the Hatano-Nelson model
where $J$ is a hopping term to the first-neighbour, $\gamma\in\mathbb{R}$ parameterised the left-right imbalance in charge hopping, also called non-reciprocity, and $c^{\dagger}_{\ell}\left(c_{\ell}\right)$ is the creation (annihilation) operator which creates (destroys) a fermion on site $\ell$. Under open boundary conditions (OBC), this non-reciprocal hopping gives rise to a unique phenomenon of non-Hermitian systems, the skin effect [67, 68, 36, 69]. This corresponds to the localization of the single-particle eigenstates at the edges of the system. In addition, the model has a huge sensitivity to the boundary conditions: under periodic boundary conditions (PBC), the single-particle spectrum encircles $E=0$ in the complex plane, with the eigenstates manifesting as delocalised plane waves. In contrast, with OBC, the spectrum is real for $\left|\gamma\right|<J$ and purely imaginary for $\left|\gamma\right|>J$. Additionally, all right single-particle eigenstates show an exponential localisation at the left (right) boundary for positive (negative) $\gamma$ (check Appendix. B for further details).
Figure 2: Left panel: Comparison between the results obtained with the Faber polynomial method and those reported in [34] for the entanglement entropy of half of the chain for a total chain of size $L=100$. Right panel: Entanglement entropy for half of the chain with $\gamma=-0.05J$ using a different number of polynomials and the time step of $0.1J^{-1}$. We note that we use a symmetric definition for $\gamma$ with regard to [34]
@@ 4.2 Domain Wall Melting for Hatano-Nelson
We start considering the time-evolution of the particle density profile under the HN dynamics, that we plot in Fig. 4 for increasing values of the non-reciprocal coupling $\gamma$. In the unitary case ($\gamma=0$) a clear light cone is visible, corresponding to ballistically propagating quasiparticles. In the non-Hermitian case a light cone is still visible at short times, at least for $\left|\gamma\right|<1$, when the particle density satisfies a scaling function $n\left(\ell/t\right)=f\left(\ell/(v_{\text{eff}}\;t\right)$, where $v_{\rm eff}$ is an effective velocity. As the non-reciprocal coupling increases, the light cone shrinks more and more, up to $\gamma=1$, corresponding to the exceptional point of the HN, at which the domain-wall state remains stable. We now focus on the short-time dynamics and discuss the origin of the velocity renormalisation. From the numerical data we obtain (see Fig. 5 (top left))
@@
We can provide a simple argument to justify the renormalisation of the velocity due to non-hermiticity. The renormalised velocity can be obtained by using a local continuity equation for the particle density (Eq. 23) by expanding the term $\mathcal{T}_{\ell}$,
In the non-interacting limit, this term can be further expanded using Wick’s theorem in a local and non-local term,
@@ 5 Application to Non-Hermitian Many-Body Systems
Figure 8: Long-time behaviour of the magnetisation profile shown in Fig (7), $\langle S_{\ell}^{z}\rangle_{\infty}$, for different values of $\Delta$. We see that at $\Delta=0$ an emergent domain wall is formed, which is stable for small $\Delta$ (left panel). Increasing $\Delta$ the system develops a potential drop in the middle of the chain, akin to diffusive dynamics, which then further develops an oscillating patter as $\Delta$ increases (right panel). Other parameters $\gamma=0.8J$.
@@ 6.1 Magnetisation and Entanglement Dynamics in Monitored Spin Chains
with =$\ell=\left\{0,\cdots,L-2\right\}$. Previous studies have shown that these operators induce a phenomenon known as the Liouvillian skin effect [99, 100, 101]. Specifically, there is an exponential localisation of the Liouvillian modes at the boundaries of the system. This becomes evident when the Linblad is projected onto the one-particle sector, where, at $J=0$ and $\Delta=0$, it simplifies to an effective Hatano-Nelson Hamiltonian tuned to its exceptional point [99],
@@
The measurement apparatus effectively disentangles the system, pushing excitations towards the left boundary, and preventing the formation of long-range correlations. It is harder to completely confirm this for smaller values of $\gamma$, where the unitary evolution dominates both the non-Hermitian and stochastic terms. This mainly affects systems with a smaller dimension, where finite-size effects are substantial, since the terms proportional to $L^{-n}$ with $n>0$ in the entanglement entropy cannot be ignored. For example, for $\gamma=0.1J$, one cannot extrapolate the true entanglement phase, as the observed linear growth might be an artefact of small system sizes. Nevertheless, the remaining data points clearly indicate the collapse of all system sizes to the same value, thus revealing the area law nature of the entanglement entropy.
@@ 7 Conclusions
This study offers additional insights into the entanglement phase transition in quantum spin chains exhibiting Non-Hermitian or Liouvillian skin effects. In the Hatano-Nelson model with quantum jumps, we found that the area law behaviour of entanglement entropy remains for any nonzero $\gamma$, similar to the no-click limit. However, the sources are different: in the no-click limit, the area law stems from the single-particle skin effect, while in the monitored stochastic trajectory, it is due to the quantum jumps that relax the system into the fully polarised down spin-state, a product state. On the other hand, the dynamics with two-body jump operators, $L_{\ell}\propto S^{+}_{\ell}S^{-}_{\ell+1}$, allow for a non-equilibrium steady state with a magnetisation profile resembling the no-click limit of the Hatano-Nelson model. Furthermore, the average entanglement entropy still follows an area law for finite values of the ratio $\gamma/J$ in conditional dynamics. The measurement apparatuses effectively disentangle the state, suppressing the volume law that would otherwise be generated in unitary dynamics. This work extends previous studies that focused on the one-particle sector [99, 34].
@@ Appendix B General Features on the Hatano-Nelson Model
with $\theta\in\mathbb{R}$ and both $p_{\ell}$ and $q_{\ell}$ two fermionic operators satisfying the unsual anticommutation relations: $\left\{p^{\dagger}_{\ell},q_{n}\right\}=\delta_{\ell,n}$, $\left\{p^{\dagger}_{\ell},p_{n}\right\}=e^{2\ell\theta}\delta_{\ell,n}$, $\left\{q^{\dagger}_{\ell},q_{n}\right\}=e^{-2\ell\theta}\delta_{\ell,n}$ and $\left\{p_{\ell},q_{n}\right\}=\left\{p^{\dagger}_{\ell},q^{\dagger}_{n}\right\}=0$. This indicates the biorthogonality of the Hamiltonian’s eigenbasis.
@@
Figure 13: Comparison between the Faber polynomial expansion and the TEBD (MPO) algorithm, shown on the left (right). The calculations on the left were performed with $\gamma=0J$, $\Delta=0.5J$ and $L=18$, while on the right we used $\gamma=0.2J$, $\Delta=0.4J$ and $L=18$.

### v2

@@ 1 Introduction
Finally, as a last example of non-unitary Markovian dynamics, we can consider the one generated by a purely non-Hermitian Hamiltonian. Non-Hermitian physics emerges intrinsically in various domains of physics beyond quantum physics, encompassing photonics [22], hydrodynamics [23] and active matter [24, 25]. In the context of open quantum systems, a non-Hermitian evolution can be obtained by post-selecting the quantum trajectories corresponding to no quantum jumps, i.e. the no-click limit [26]. Non-Hermitian quantum systems show anomalous static and dynamical properties, which are attracting widespread interest. Among these, we mention the unconventional propagation of quantum correlations [27, 28, 29], the distinct entanglement transitions generated by time evolution [30, 31, 32, 33, 34] or their extraordinary sensitivity to boundary conditions, also known as the skin effect [35, 36, 37, 38, 39, 40]. The latter manifests itself through the unusual localisation of all single-particle eigenstates at the system’s edges under open boundary conditions  [41, 42, 43, 44, 45]. Furthermore, it has unique signatures in the dynamics, leading to non-reciprocal transport[46].
@@ 3 Faber Polynomial Method
where we define |$\ket{\Psi_{n}}=\mathcal{P}_{n}\left(\mathcal{H}\right)\ket{\Psi\left(t_{0}\right)}$. As will be demonstrated, the coefficients $c_{n}(\delta t)$ decrease as the order $n$ increases. Moreover, the states |$\ket{\Psi_{n}}$ are efficiently computed through the recurrence relations that the polynomials satisfy. To compute the subsequent level of the expansion, the main computational task is the application of the system’s Hamiltonian onto a particular state. Consequently, the most demanding operation involves only multiplying the Hamiltonian by a limited group of vectors, leading to a resource usage that increases linearly with $\dim\left(\mathcal{H}\right)$ for sparse matrices or quadratically with $\dim\left(\mathcal{H}\right)$ for dense matrices. Linear scaling is expected for Hamiltonians characterising a system with short-range interactions or hoppings. These principles are exactly those underpinning Kernel Polynomial Methods [64], which has become an essential computational resource in condensed matter physics, particularly for calculating various spectral quantities[65, 66, 67, 68, 69].
@@ 3.2 Non-Unitary Evolution
In order to perform one time step of evolution, one simply has to truncate Eq. 10 up to the desired order $N_{p}$, and calculate the associated states |$\ket{\Psi_{n}}$. Through the relations in Eq. 12, one never has to store the Hamiltonian matrix, only needing to store in memory at most the previous two states, |$\ket{\Psi_{n-1}}$ and |$\ket{\Psi_{n}}$, to compute the following term of the expansion |$\ket{\Psi_{n+1}}$. Furthermore, the expansion coefficients can be computed once at the beginning of the algorithm, as they depend only on the chosen time step. Given this, the computation of the state of the system after a time step scales linearly in the number of polynomials and linearly in the Hilbert space dimension (assuming that the Hamiltonian has a sparse representation on the used basis). This scaling can be improved by using parallelisation techniques [68] and making use of the underlying symmetries of the Hamiltonian [72].
@@ 4.1 Benchmark: Dynamics of the Hatano-Nelson model
where $J$ is a hopping term to the first-neighbour, $\gamma\in\mathbb{R}$ parameterised the left-right imbalance in charge hopping, also called non-reciprocity, and $c^{\dagger}_{\ell}\left(c_{\ell}\right)$ is the creation (annihilation) operator which creates (destroys) a fermion on site $\ell$. Under open boundary conditions (OBC), this non-reciprocal hopping gives rise to a unique phenomenon of non-Hermitian systems, the skin effect [35, 36, 37, 38, 39, 40]. This corresponds to the localization of the single-particle eigenstates at the edges of the system. In addition, the model has a huge sensitivity to the boundary conditions: under periodic boundary conditions (PBC), the single-particle spectrum encircles $E=0$ in the complex plane, with the eigenstates manifesting as delocalised plane waves. In contrast, with OBC, the spectrum is real for $\left|\gamma\right|<J$ and purely imaginary for $\left|\gamma\right|>J$. Additionally, all right single-particle eigenstates show an exponential localisation at the left (right) boundary for positive (negative) $\gamma$ (check Appendix. B for further details).
Figure 2: Left panel: Comparison between the results obtained with the Faber polynomial method and those reported in [34] for the entanglement entropy of half of the chain for a total chain of size $L=100$. Right panel: Entanglement entropy for half of the chain, with $\gamma=-0.05J$ using a different number of polynomials and the time step of $0.1J^{-1}$. We note that we use a symmetric definition for $\gamma$ with regard to [34]
@@ 4.2 Domain Wall Melting for Hatano-Nelson
We start considering the time-evolution of the particle density profile under the HN dynamics, that we plot in Fig. 4 for increasing values of the non-reciprocal coupling $\gamma$. In the unitary case ($\gamma=0$) a clear light cone is visible, corresponding to ballistically propagating quasiparticles. In the non-Hermitian case a light cone is still visible at short times, at least for $\left|\gamma\right|<J$, when the particle density satisfies a scaling function $n\left(\ell/t\right)=f\left(\ell/(v_{\text{eff}}\;t\right)$, where $v_{\rm eff}$ is an effective velocity. As the non-reciprocal coupling increases, the light cone shrinks more and more, up to $\gamma=J$, corresponding to the exceptional point of the HN, at which the domain-wall state remains stable. We now focus on the short-time dynamics and discuss the origin of the velocity renormalisation. From the numerical data we obtain (see Fig. 5 (top left))
@@
We can provide a simple argument to justify the renormalisation of the velocity due to non-hermiticity. The renormalised velocity can be obtained by using a local continuity equation for the particle density (Eq. 23) by working the term $\mathcal{T}_{\ell}$. In the non-interacting limit, this term can be expanded using Wick’s theorem in a local and non-local term,
@@ 5 Application to Non-Hermitian Many-Body Systems
Figure 8: Long-time behaviour of the magnetisation profile shown in Fig. 7, $\langle S_{\ell}^{z}\rangle_{\infty}$, for different values of $\Delta$. We see that at $\Delta=0$ an emergent domain wall is formed, which is stable for small $\Delta$ (left panel). Increasing $\Delta$ the system develops a potential drop in the middle of the chain, akin to diffusive dynamics, which then further develops an oscillating patter as $\Delta$ increases (right panel). Other parameters $\gamma=0.8J$.
@@ 6.1 Magnetisation and Entanglement Dynamics in Monitored Spin Chains
with =$\ell=\left\{0,\cdots,L-2\right\}$. Previous studies have shown that these operators induce a phenomenon known as the Liouvillian skin effect [24, 105, 106, 107]. Specifically, there is an exponential localisation of the Liouvillian modes at the boundaries of the system. This becomes evident when the Linblad is projected onto the one-particle sector, where, at $J=0$ and $\Delta=0$, it simplifies to an effective Hatano-Nelson Hamiltonian tuned to its exceptional point [105],
@@
The measurement apparatus effectively disentangles the system, pushing excitations towards the left boundary, and preventing the formation of long-range correlations. It is harder to completely confirm this for smaller values of $\gamma$, where the unitary evolution dominates both the non-Hermitian and stochastic terms. This mainly affects systems with a smaller dimension, where finite-size effects are substantial, since the terms proportional to $L^{-n}$ with $n>0$ in the entanglement entropy cannot be ignored. For example, for $\gamma=0.1J$, one cannot extrapolate the true entanglement scaling, as the observed linear growth might be an artefact of small system sizes. Nevertheless, the remaining data points clearly indicate the collapse of all system sizes to the same value, thus revealing the area law nature of the entanglement entropy.
@@ 7 Conclusions
This study offers additional insights into the entanglement transition in quantum spin chains exhibiting Non-Hermitian or Liouvillian skin effects. In the Hatano-Nelson model with quantum jumps, we found that the area law behaviour of entanglement entropy remains for any nonzero $\gamma$, similar to the no-click limit. However, the sources are different: in the no-click limit, the area law stems from the single-particle skin effect, while in the monitored stochastic trajectory, it is due to the quantum jumps that relax the system into the fully polarised down spin-state, a product state. On the other hand, the dynamics with two-body jump operators, $L_{\ell}\propto S^{+}_{\ell}S^{-}_{\ell+1}$, allow for a non-equilibrium steady state with a magnetisation profile resembling the no-click limit of the Hatano-Nelson model. Furthermore, the average entanglement entropy still follows an area law for finite values of the ratio $\gamma/J$ in conditional dynamics. The measurement apparatuses effectively disentangle the state, suppressing the volume law that would otherwise be generated in unitary dynamics. This work extends previous studies that focused on the one-particle sector [105, 34].
@@ Appendix B General Features on the Hatano-Nelson Model
with $\theta\in\mathbb{R}$ and both $p_{\ell}$ and $q_{\ell}$ two fermionic operators satisfying the unusual anticommutation relations: $\left\{p^{\dagger}_{\ell},q_{n}\right\}=\delta_{\ell,n}$, $\left\{p^{\dagger}_{\ell},p_{n}\right\}=e^{2\ell\theta}\delta_{\ell,n}$, $\left\{q^{\dagger}_{\ell},q_{n}\right\}=e^{-2\ell\theta}\delta_{\ell,n}$ and $\left\{p_{\ell},q_{n}\right\}=\left\{p^{\dagger}_{\ell},q^{\dagger}_{n}\right\}=0$. This indicates the biorthogonality of the Hamiltonian’s eigenbasis.
@@
Figure 13: Comparison between the Faber polynomial expansion and the TEBD (MPO) algorithm, shown on the left (right). The calculations on the left were performed with $\gamma=0.0J$, $\Delta=0.5J$ and $L=18$, while on the right we used $\gamma=0.2J$, $\Delta=0.4J$ and $L=18$.